# AI Agentic Programming: A Survey of Techniques, Challenges, and Opportunities

### v1

@@ 3. Survey Methodology
This survey follows a widely-used systematic literature review (SLR) methodology (keele2007guidelines, ; DBLP:journals/corr/Gong24Deep, ; hou2024largelanguagemodelssoftware, ; DBLP:journals/tse/WangHGGZFSLZN23, ; zhang2024systematic, ; qudus2025fact, ) to provide comprehensive coverage of AI agentic programming research, as illustrated in Figure 5.
@@ 3.1. Search Strategy
We conducted automatic searches across multiple academic databases, including Google Scholar, ACM Digital Library, IEEE Xplore, SpringerLink, and arXiv.org. We also examined proceedings from top-tier venues (FSE, ICSE, ASE, ICML, NeurIPS, AAAI, etc.).
@@ 3.2. Study Selection
After initial retrieval, we followed a three-stage study selection process: (1) title and abstract screening by two independent researchers, (2) full-text review with disagreement resolution through discussion, and (3) backward and forward citation chaining to identify additional relevant studies. During the selection process, we used the following criteria:
@@ 3.3. Results
Among the 152 academic references published from 2022 to 2025 (excluding tool descriptions and websites), 5% appeared in 2022, 22% in 2023, 53% in 2024, and 20% in 2025, as shown in Figure 5, reflecting a surge in AI agent programming research following the widespread adoption of LLMs.
Figure 4. Survey Methodology.
@@
[width=]figures/pie_year
Figure 5. Distribution of Papers.
@@ 4.1. Agentic Behaviour Dimensions
Single-turn agents perform actions in response to individual prompts, often without preserving context. For example, classical GitHub Copilot responds to each prompt independently, without remembering past interactions. In contrast, multi-turn agents, such as GitHub Copilot Agent or Claude Opus 4 with tooling capabilities, maintain state across interactions, enabling iterative refinement, exploration, and goal pursuit. For instance, GitHub Copilot Agent can hold a conversation across multiple steps, remember earlier function names, and build a complete module through back-and-forth iterations between agents (github_copilot, ).
@@
Static agents (e.g., GitHub Copilot and Tabnine) follow predefined workflows or heuristics. Adaptive agents modify their strategies using feedback from tools, user input, or environmental signals. Some employ learning mechanisms to improve over time. For example, GitHub Copilot Agent adapts its approach when test failures occur, revising its implementation or replanning subtasks.
@@ 4.2.1. Interactive Code Assistants
GitHub Copilot (github_copilot, ) and Cursor (cursor, ) are two representative examples of LLM-based code assistants. GitHub Copilot, originally developed based on Codex trained on GitHub code repositories, offers context-aware code completions across multiple programming languages and is tightly integrated into popular IDEs such as Visual Studio Code (vscode, ) and JetBrains. Cursor extends this functionality by embedding conversational interaction, maintaining memory of previous edits, and supporting structured command execution, allowing for a more dynamic and iterative user experience. Other notable systems include Amazon Q Develiper (amazon-q-developer, ) and Tabnine (tabnine, ). Amazon CodeWhisperer targets developers working in cloud ecosystems, offering language-specific completions, cloud API integration, and basic vulnerability detection. In contrast to CodeWhisperer, Tabnine takes a privacy-first approach by deploying smaller local models trained on permissively licensed code, making it attractive for organizations and developers who do not want to send their code to a remote cloud.
@@ 4.2.2. Autonomous Task-Oriented Agents
These agents perform multi-step programming tasks with minimal human intervention, often maintaining control over the entire development process from requirement interpretation to code generation and validation. They can plan, execute, and revise their own workflows in response to intermediate results or changing task requirements. Many integrate external tools such as debuggers, package managers, or search engines, enabling them to gather information, resolve errors, and optimize code without direct user guidance. Examples include the state-of-the-art LLMs that have tool integration capabilities like GPT-5 (openai2025gpt5introducing, ), Claude Opus 4 (anthropic_opus4, ), Google Jules (Jules, ), Kimi K2 (team2025kimi, ), and Gemini 2 (comanici2025gemini25pushingfrontier, ). These agents are often proactive in suggesting next steps, adaptive to new inputs, and capable of maintaining continuity across extended sessions through persistent memory or context management mechanisms.
@@ 4.2.3. Planning-Centric Agents
Planning-centric agents approach problem-solving as a two-phase process: first, structured task decomposition, where high-level goals are broken down into smaller, more manageable steps; and second, execution monitoring, in which results are evaluated and the plan is adjusted accordingly. This approach improves the handling of long-horizon tasks. For example, CAMEL (li2023camel, ) employs two agents in a role-playing setup, a “user” agent and an “assistant” agent, which collaboratively refine goals and strategies, producing plans that downstream code generators can execute. These agents are typically multi-turn, memory-enabled, and trading speed for robustness.
@@ 4.2.4. Multi-Agent and Collaborative Systems
Multi-agent and collaborative systems extend agent-based programming by introducing multiple specialized agents that coordinate to solve complex software engineering tasks. This approach draws inspiration from human software teams, where each member has a distinct role, such as requirements analysis, coding, testing, or documentation, and communication protocols are established to ensure progress toward shared objectives. For example, SWE-Agent (yang2024swe, ) employs multiple role-specific LLM agents—an “Architect” agent for high-level design, a “Coder” agent for implementation, and a “Reviewer” agent for quality assurance—connected through structured dialogue and shared memory. Such systems offer high autonomy, scalability, and tool integration, but face challenges in avoiding redundancy and managing conflicts.
@@ 4.3. Summary and Comparative Table
GitHub Copilot IDE Assistant Reactive
Tabnine IDE Assistant Reactive
Cursor IDE Assistant Reactive
Claude Opus 4 Task-oriented Proactive
Google Jules Task-oriented Proactive
Dev-GPT Task-oriented Proactive
KimI K2 Task-oriented Proactive
Gemini 2 Task-oriented Proactive
Voyager Planning Agent Proactive
CAMEL Planning Agent Proactive
CodePlan Planning Agent Proactive
ChatDev Multi-agent System Proactive
AutoCodeRover Multi-agent System Proactive
SWE-Agent Multi-agent System Proactive
@@ 4.4. Cost and Token Consumption Model
Table 5. Representative token pricing for LLMs used in coding tasks (USD per 1M tokens, as of Aug. 2025).
@@
Kimi K2 0.15 2.50 128k
Qwen3-235B-A22B 0.22 0.88 32k
Qwen3-Coder-480B-A35B-Instruct 4.5 13.5 256k
Solar-Pro 0.30 0.30 128k
@@
Devstral-Small (Mistral) 0.10 0.30 128k
Devstral-Large (Mistral Medium 3) 0.40 2.00 128k
@@ 5. Challenges
AI agentic programming introduces a promising and complex shift in how software is developed, relying on the autonomous capabilities of LLMs. Despite recent progress, several technical and conceptual challenges remain that hinder the deployment of robust, scalable, and trustworthy agentic systems (sapkota2025aiagentsvsagentic, ; acharya2025agentic, ).
5.1. Inadequate Evaluation and Benchmarking
@@
A variety of benchmarks and open-source toolkits (liu2023agentbench, ; qin2023toolllmfacilitatinglargelanguage, ) have been proposed to evaluate the capabilities of LLM-based agents across different tasks. Despite this progress, existing benchmarks commonly used for assessing coding agents, such as HumanEval (zheng2023codegeex, ) and SWE-Bench (jimenez2023swe, ), remain inadequate for capturing the full complexity of real-world software engineering workflows.
@@
These benchmarks primarily focus on small, self-contained problems, often restricted to a narrow range of programming languages, such as Python (chen2021evaluatinglargelanguagemodels, ), and generally lack support for interactive, multi-turn, or tool-integrated tasks (liu2023agentbench, ). In contrast, practical agentic systems are expected to operate over large, modular codebases, interface with third-party libraries, manage build workflow pipelines, and respond dynamically to user feedback or runtime tool outputs. As LLM-based systems increasingly incorporate reinforcement learning (ouyang2022training, ; wang2024survey, ) and more advanced planning mechanisms, future benchmarks should reflect this integration.
@@
Moreover, there is a noticeable absence of evaluation frameworks designed for emerging complex use cases, such as those involving interactions with compilers and debuggers (wang2024executablecodeactionselicit, ), where agents must reason about low-level program behavior, perform iterative transformations, or track state across toolchains. The lack of such domain-specific benchmarks presents a significant gap in evaluating agent performance under realistic conditions.
5.2. Domain Foundation Models for Agents
Generic coding agents often struggle in domain-specific environments such as embedded systems, high-performance computing, optimization, or formal software verification (lecong2025llmsreasonprogramsemantics, ; loughridge2024dafnybench, ). These domains typically impose stricter operational constraints and require deep integration with specialized APIs, toolchains, and domain knowledge resources that are often underrepresented in general-purpose training corpora.
@@ 5.2. Domain Foundation Models for Agents
To address these limitations, recent research has proposed domain-adapted models and task-specific learning strategies to accelerate agent performance in specialized settings (song2025injecting, ). For instance, some approaches have begun incorporating compiler knowledge or security-specific patterns into LLM training pipelines (lin2025llmssecuritycoursesecuring, ; LLMCompiler, ), enabling agents to reason more effectively about low-level program behavior or vulnerability patterns.
In the future, developing robust and adaptable domain foundation models will be a promising direction for enabling agents to operate reliably in complex software environments, such as LLMs pretrained or fine-tuned on domain-specific data, tools, and semantics.
@@ 5.3. Safety and Privacy
As agentic systems gain increasing autonomy, so does the potential for unsafe behavior. Unlike traditional tools (Checkmarx, ), agentic systems can invoke external tools, perform structural code modifications, and even commit changes without direct human oversight (triedman2025multiagentsystemsexecutearbitrary, ; zhang2025agentsafetybenchevaluatingsafetyllm, ). These capabilities introduce significant risks, including the possibility of introducing subtle bugs, propagating unsafe patterns, or violating security constraints.
A critical future direction involves ensuring that agentic systems can protect users and data. For example, when agents visit private repositories or are deployed in cloud-integrated environments, future models may need built-in controls to restrict access to sensitive project data.
Further, malicious prompts, poisoned APIs, or compromised toolchains can mislead agents into executing unsafe behaviors (debenedetti2024agentdojodynamicenvironmentevaluate, ; zhan2024injecagentbenchmarkingindirectprompt, ). Future research should prioritize the design of secure protocols for agent collaboration, including authentication between agents, validation of tool outputs, and detection of anomalous actions.
Also, agents must be capable of explaining their reasoning, flagging uncertainties, and allowing developers to understand and revise with minimal effort. Building safety and privacy into the foundation of agentic architectures is essential.
@@ 5.4. Toolchain Integration and Programming Language Design
One fundamental challenge lies in the incompatibility between existing software tools and the needs of autonomous agents. Most programming languages, compilers, debuggers, and development tools are designed with human developers in mind (rorseth2025ladybug, ).
They emphasize usability over feedback, so agents often struggle to diagnose failures, trace the effects of their code transformations, or understand build errors (kim2024llm, ). For example, compilers typically provide minimal insights into transformation failures or semantic conflicts (li2025review, ).
To support agentic workflows, toolchains must evolve to expose IRs, transformation traces, and structured feedback interfaces. Moreover, language-level annotations and agent-aware interfaces are needed to communicate developer intent and guide automated reasoning.
@@ 5.5. Scalable Memory
Agentic programming systems must maintain coherence and reasoning over long-running tasks involving multiple iterations, tools, and contextual dependencies. However, current LLMs are limited by fixed context windows and lack persistent, structured memory mechanisms (wang2023augmenting, ).
Realistic software tasks, such as feature implementation, debugging, or refactoring, require agents to store and reason over evolving states, feedback logs, intermediate plans, and prior actions (qian2025memorag, ). Without hierarchical and queryable memory systems, agents risk repeating errors, forgetting past successes, or producing inconsistent results.
Emerging solutions such as retrieval-augmented generation and memory summarization offer partial relief, but they remain inadequate for complex, multi-session workflows (maharana2024evaluating, ). Future research can explore memory architectures that differentiate short-term interactions, mid-term subgoals, and long-term domain knowledge.

### v2

@@ 3. Survey Methodology
This survey follows a widely-used systematic literature review (SLR) methodology (Kitchenham and Charters, 2007; Gong and Chen, 2024a; Hou et al., 2024; Wang et al., 2023b; Zhang et al., 2024a; Qudus et al., 2025) to provide comprehensive coverage of AI agentic programming research, as illustrated in Figure 6.
Figure 5. Survey methodology for academic research.
[…]
[width=]figures/pie_year
Figure 6. Distribution of academic papers.
@@ 3.1. Search Strategy
We conducted automatic searches across multiple academic databases, including Google Scholar, ACM Digital Library, IEEE Xplore, SpringerLink, and arXiv.org. We also examined proceedings from top-tier venues (FSE, ICSE, ASE, ICML, NeurIPS, AAAI, etc.). As this fast-evolving field is largely dominated by industry, we also pay attention to open-source and the industry releases of relevant results and tools.
@@ 3.2. Study Selection
After initial retrieval, we followed a three-stage study selection process for academic papers: (1) title and abstract screening by two independent researchers, (2) full-text review with disagreement resolution through discussion, and (3) backward and forward citation chaining to identify additional relevant studies. During the selection process, we used the following criteria:
@@ 3.3. Results
  • Software tools and industry products: we also study a wide range of state-of-the-art AI coding agents and LLMs like GitHub Copilot Agents, GPT, Gemini, Deepseek, Qwen and Claude Opus 4.
Among the 152 academic references published from 2022 to 2025 (excluding tool descriptions and websites), 5% appeared in 2022, 22% in 2023, 53% in 2024, and 20% in 2025, as shown in Figure 6, reflecting a surge in AI agent programming research following the widespread adoption of LLMs. Based on our systematic analysis, we developed a hierarchical classification of AI agentic programming systems along behavioral characteristics and system architectures, as shown in Figure 7. The following section examines each category in detail.
@@
[width=]figures/tree
Figure 7. Taxonomy of AI agentic programming systems.
@@ 4.1. Agentic Behaviour Dimensions
Single-turn agents perform actions in response to individual prompts, often without preserving context. For example, classical GitHub Copilot responds to each prompt independently, without remembering past interactions. In contrast, multi-turn agents, such as GitHub Copilot Agent or Claude Opus 4 with tooling capabilities, maintain state across interactions, enabling iterative refinement, exploration, and goal pursuit. For instance, GitHub Copilot Agent can hold a conversation across multiple steps, remember earlier function names, and build a complete module through back-and-forth iterations between agents (GitHub, 2021).
@@
Static agents (e.g., GitHub Copilot and Tabnine (Tabnine, 2025)) follow predefined workflows or heuristics. Adaptive agents modify their strategies using feedback from tools, user input, or environmental signals. Some employ learning mechanisms to improve over time. For example, GitHub Copilot Agent adapts its approach when test failures occur, revising its implementation or replanning subtasks.
@@ 4.2.1. Interactive code assistants
GitHub Copilot (GitHub, 2021) and Cursor (cur, 2025) are two representative examples of LLM-based code assistants. GitHub Copilot, originally developed based on Codex trained on GitHub code repositories, offers context-aware code completions across multiple programming languages and is tightly integrated into popular IDEs like Visual Studio Code (vsc, 2025) and JetBrains. Cursor extends this functionality by embedding conversational interaction, maintaining memory of previous edits, and supporting structured command execution. Other notable systems include Amazon Q Developer (Amazon Web Services, 2025) and Tabnine (Tabnine, 2025). Amazon Q Developer targets developers working in cloud ecosystems, offering language-specific completions, cloud API integration, and basic vulnerability detection. In contrast to Amazon Q Developer, Tabnine takes a privacy-first approach by deploying smaller local models trained on permissively licensed code, making it attractive for organizations and developers who do not want to send their code to a remote cloud.
@@ 4.2.2. Autonomous task-oriented agents
These agents perform multi-step programming tasks with minimal human intervention, often maintaining control over the entire development process from requirement interpretation to code generation and validation. They can plan, execute, and revise their own workflows in response to intermediate results or changing task requirements. Many integrate external tools such as debuggers, package managers, or search engines, enabling them to gather information, resolve errors, and optimize code without direct user guidance. Examples include GPT-5 (OpenAI, 2025), which functions as an end-to-end collaborator for long-horizon code generation and debugging; Claude Opus 4 (Anthropic, 2025b), engineered for sustained agentic workflows with strong long-context and tool-integration capabilities; Google Jules (Jul, 2025), which can sandbox repositories, propose diffs, and execute verified changes on real projects; DevGPT (devgpt-labs, 2024), an open-source pipeline that converts tickets into actionable code and CI artifacts; Kimi K2 (Team et al., 2025), a model optimized for agentic coding proficiency; and Gemini 2 (et al., 2025), which emphasizes multimodal inputs and built-in planning modes. These agents are often proactive in suggesting next steps, adaptive to new inputs, and capable of maintaining continuity across extended sessions through persistent memory or context management mechanisms.
@@ 4.2.3. Planning-centric agents
Planning-centric agents approach problem-solving as a two-phase process: first, structured task decomposition, where high-level goals are broken down into smaller, more manageable steps; and second, execution monitoring, in which results are evaluated and the plan is adjusted accordingly. This approach improves the handling of long-horizon tasks. For example, CAMEL (Li et al., 2023b) employs two agents in a role-playing setup, a “user” agent and an “assistant” agent, which collaboratively refine goals and strategies, producing plans that downstream code generators can execute. Voyager (Wang et al., 2023c) demonstrates this paradigm by continuously exploring an open-ended environment, generating executable skills, and refining them into reusable plans for long-term adaptability. Similarly, CodePlan (Wen et al., 2024) introduces code-form planning, where structured pseudocode serves as an explicit intermediate representation to decompose complex problems into executable steps. These agents are typically multi-turn, memory-enabled, and trading speed for robustness.
@@ 4.2.4. Multi-agent and collaborative systems
Multi-agent and collaborative systems extend agent-based programming by introducing multiple specialized agents that coordinate to solve complex software engineering tasks. This approach draws inspiration from human software teams, where each member has a distinct role, such as requirements analysis, coding, testing, or documentation, and communication protocols are established to ensure progress toward shared objectives. For example, SWE-Agent (Yang et al., 2024b) employs multiple role-specific LLM agents: an “Architect” agent for high-level design, a “Coder” agent for implementation, and a “Reviewer” agent for quality assurance, which are connected through structured dialogue and shared memory. ChatDev (Qian et al., 2024) follows this paradigm by simulating an end-to-end software company, where agents take on roles such as CEO, CTO, and programmers to collaboratively design, implement, and test applications. Furthermore, AutoCodeRover (Zhang et al., 2024e) extends collaboration into real-world repositories, orchestrating specialized agents that autonomously navigate, edit, and validate source code across complex multi-file projects.
@@ 4.3. Summary and Comparative Table
[…]
@@ 4.4. Cost and Token Consumption Model
Table 6. Example token pricing for LLMs used in coding tasks (USD per 1M tokens, as of Sep. 2025).
@@
[…]
Qwen3-235B-A22B 0.22 0.88 128k
[…]
@@
[…]
Devstral-Medium 0.40 2.00 128k
@@ 5. Challenges
AI agentic programming introduces a promising and complex shift in how software is developed, relying on the autonomous capabilities of LLMs. Despite recent progress, several technical and conceptual challenges remain that hinder the deployment of robust, scalable, and trustworthy agentic systems (Sapkota et al., 2025; Acharya et al., 2025).
5.1. Evaluation and Benchmarking
@@
A variety of benchmarks and open-source toolkits (Liu et al., 2023e; Qin et al., 2023) have been proposed to evaluate the capabilities of LLM-based agents across different tasks. Despite this progress, existing benchmarks commonly used for assessing coding agents, such as HumanEval (Zheng et al., 2023b) and SWE-Bench (Jimenez et al., 2023), may still be inadequate for capturing the full complexity of real-world software engineering workflows.
@@
Most of these benchmarks primarily focus on small, self-contained problems, often restricted to a small number of programming languages, such as Python (Chen et al., 2021a), and generally lack support for interactive, multi-turn, or tool-integrated tasks (Liu et al., 2023e). In contrast, practical agentic systems are expected to operate over large, modular codebases, interface with third-party libraries, manage build workflow pipelines, and respond dynamically to user feedback or runtime tool outputs. As LLM-based systems increasingly incorporate reinforcement learning (Ouyang et al., 2022; Wang et al., 2024e) and more advanced planning mechanisms, future benchmarks should reflect this integration. Table 7 shows the characteristics of tasks in SWE-Bench. Although SWE-Bench introduces project-level repositories and leverages unit tests and continuous integration (CI) for evaluation, its scope is restricted to Python. Most tasks are function- or module-level, with no support for multi-turn feedback, third-party library usage, or build pipeline management. Even for project-level tasks, interaction is limited to binary pass/fail outcomes, providing only minimal support for realistic multi-step or tool-integrated software engineering workflows.
@@
Moreover, there is a noticeable absence of evaluation frameworks designed for emerging complex use cases, such as those involving interactions with compilers and debuggers (Wang et al., 2024b), where agents must reason about low-level program behavior, perform iterative transformations, or track state across toolchains. The lack of such domain-specific benchmarks presents a significant gap in evaluating agent performance under realistic conditions.
Table 7. SWE-Bench characteristics.
Task type Proportion Languages Interaction Multi-turn feedback Library integration Build pipeline
Function-level  65% Python Unit tests
[…]
Project-level ¡10% Python Tests and CI Limited (pass/fail)
5.2. Communication Protocols for Multi-Agent Systems
Early protocols enabled one-to-one agent–to–tool interactions but did not support direct agent-to-agent communication (MCP, 2025). Current practice often relies on heterogeneous web service protocols and adapters, which provide interoperability but introduce high latency, bandwidth overhead, and limited scalability (Naik and Jenkins, 2016). Recent work (Xu et al., 2012; A2A, 2025; Verma et al., 2015) explores group session models, where each execution unit functions as a session or group agent coordinating multiple services, and standards efforts propose unified message formats and session semantics. However, the lack of a common protocol, inefficiencies of adapters, and challenges in managing dynamic multi-agent sessions remain open problems for scalable and dependable agentic systems.
5.3. Domain-specific Models for Agents
@@
Generic coding agents often struggle in domain-specific environments such as embedded systems, high-performance computing, optimization, or formal software verification (Le-Cong et al., 2025; Loughridge et al., 2024). These domains typically impose stricter operational constraints and require deep integration with specialized APIs, toolchains, and domain knowledge resources that are often underrepresented in general-purpose training corpora. To address these limitations, recent research has proposed domain-adapted models and task-specific learning strategies to accelerate agent performance in specialized settings (Song et al., 2025). For instance, some approaches have begun incorporating compiler knowledge or security-specific patterns into LLM training pipelines (Lin et al., 2025; Cummins et al., 2025), enabling agents to reason more effectively about low-level program behavior or vulnerability patterns. In the future, developing robust and adaptable domain foundation models will be a promising direction for enabling agents to operate reliably in complex software environments, such as LLMs pretrained or fine-tuned on domain-specific data, tools, and semantics.
@@ 5.4. Safety and Privacy
As agentic systems gain increasing autonomy, so does the potential for unsafe behavior. Unlike traditional tools (Che, 2025), agentic systems can invoke external tools, perform structural code modifications, and even commit changes without direct human oversight (Triedman et al., 2025; Zhang et al., 2025). These capabilities introduce significant risks, including the possibility of introducing subtle bugs, propagating unsafe patterns, or violating security constraints. A critical future direction involves ensuring that agentic systems can protect users and data. For example, when agents visit private repositories or are deployed in cloud-integrated environments, future models may need built-in controls to restrict access to sensitive project data. Further, malicious prompts, poisoned APIs, or compromised toolchains can mislead agents into executing unsafe behaviors (Debenedetti et al., 2024; Zhan et al., 2024). Future research should prioritize the design of secure protocols for agent collaboration, including authentication between agents, validation of tool outputs, and detection of anomalous actions. Also, agents must be capable of explaining their reasoning, flagging uncertainties, and allowing developers to understand and revise with minimal effort. Building safety and privacy into the foundation of agentic architectures is essential.
@@ 5.5. Toolchain Integration and Programming Language Design
One fundamental challenge is the incompatibility between existing software tools and the needs of autonomous coding agents. Most programming languages, compilers, debuggers, and development environments were designed for human developers (Rorseth et al., 2025). They emphasize usability and readability over structured, machine-consumable feedback. As a result, agents often struggle to diagnose failures, trace the consequences of code transformations, or interpret build errors (Kim et al., 2024). For example, compilers typically report transformation failures or semantic conflicts in the form of coarse error messages, providing little insight into why an optimization was blocked or how a type error emerged (Li, 2025). Languages similarly prioritize human readability over machine-negotiated meaning, while compilers conceal internal reasoning to avoid overwhelming human users. These design choices, while historically effective, now limit the ability of AI coding agents to construct safe, efficient, and adaptive software.
To enable agentic workflows, toolchains should evolve to expose richer intermediate representations, transformation traces, and structured feedback interfaces. Equally important, programming languages should incorporate annotations and agent-aware interfaces that make developer intent explicit, allowing automated reasoning to be guided by semantic contracts rather than inferred heuristics. Together, these advances would transform compilers and languages from opaque tools into collaborative infrastructures capable of supporting autonomous agents at scale.
@@ 5.6. Scalable Memory
Agentic programming systems must maintain coherence and reasoning over long-running tasks involving multiple iterations, tools, and contextual dependencies. However, current LLMs are limited by fixed context windows and lack persistent, structured memory mechanisms (Wang et al., 2023a). Realistic software tasks may require agents to store and reason over evolving states, feedback logs, intermediate plans, and prior actions (Qian et al., 2025). Without hierarchical and queryable memory systems, agents risk repeating errors, forgetting past successes, or producing inconsistent results. Emerging solutions such as retrieval-augmented generation and memory summarization offer partial relief, but they remain inadequate for complex, multi-session workflows (Maharana et al., 2024). Future research can explore memory architectures that differentiate short-term interactions, mid-term subgoals, and long-term domain knowledge.